The Head of the Math Department announced that the mean score of Grade 11 students in the first

midterm examination in Statistics and Probability was 89 and the standard deviation was 12. One

student believed that the mean score was less than this, randomly selected 34 students and computed

their mean score. She obtained a mean score of 85. At 0.01 of level of significance, test the student’s


belief.


1) What are the null hypotheses and alternative hypotheses of the research scenario?

2) What is the level of significance?

3) What is the test statistic?

4) What is the critical value in the scenario?

5) What could be the conclusion?


1
Expert's answer
2022-04-04T18:03:08-0400

We have that:

"\mu=89"

"\sigma=12"

n=34

x=85

"\alpha=0.01"

"H_0: \mu=89"

"H_1 :\mu<89"

The hypothesis test is left-tailed.


Since the population standard deviation is known and the sample size is large (>30) we use the z-test.

The critical value for "\alpha=0.01" is "z_{0.01}=-2.33"

The critical region is "z<-2.33"

Test statistic:

"Z_{test}=\frac{x-\mu}{\frac{s}{\sqrt{n}}}=\frac{85-89}{\frac{12}{\sqrt{34}}}=-1.94"

Since –1.94 > –2.33 thus the Ztest does not fall in the rejection region we fail to reject the null hypothesis.


There is no sufficient evidence to support the student’s belief. We are 99% confident to conclude that the mean score of Grade 9 students in the first periodic examination in Mathematics is 89.